A particle from rest accelerates the magnitude of 3 m/s^2. What is its distance after 8 seconds of travel?


1
Expert's answer
2021-09-21T13:57:22-0400
"d=v_0t+\dfrac{1}{2}at^2,""d=0+\dfrac{1}{2}\cdot3\ \dfrac{m}{s^2}\cdot(8\ s)^2=96\ m."